OpenStudy (anonymous):

Write the equation of the line that passes through the point (6, -2) and has a slope of -2/3

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

y = -(2/3)(x-6)-2

OpenStudy (amistre64):

\[y = -\frac{2x}{3} -6\]

OpenStudy (amistre64):

that aint right is it..

OpenStudy (anonymous):

idk lol

OpenStudy (amistre64):

-*-=+ i see it

OpenStudy (amistre64):

\[y = -\frac{2}{3}x-10\] maybe

OpenStudy (amistre64):

im gonna have to do this the long way...

OpenStudy (anonymous):

y=−(2/3)X + 2

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

y = -(2/3)(x-6)-2 y = (-2/3)x+(-2/3)-6 -2 y = (-2/3)x+(12/3) -2 y = (-2/3)x+2 is right

OpenStudy (amistre64):

....fraction keep laughing at me

OpenStudy (anonymous):

you need to use the point slope formula for these problems: \[(y-y _{1}) = m(x - x _{1})\]

OpenStudy (amistre64):

i shoulda multied his butt with a 3 ....thatd a showed it

OpenStudy (anonymous):

so this isnt the right answer

OpenStudy (anonymous):

just use the formula, y-y1=m(x-x1) (6, -2) and has a slope of -2/3 y+2=-2/3(x-6) 3(y+2)=-2(x-6) 3y+6=-2x+12 2x+3y=12-6 2x+3y=6

OpenStudy (anonymous):

this is right.